OpenStudy (anonymous):

dy/dx = y/(x+1)(x+2) differential equation

OpenStudy (amistre64):

looks seperable to me

OpenStudy (anonymous):

1/y dy = 1/(x+1)(x+2)

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

integrate both sides :)

OpenStudy (amistre64):

might want to partial out the right side ... but yes

OpenStudy (anonymous):

i dont know how to integrate the right hand side...

OpenStudy (anonymous):

erm, my partial fractions is a bit rusty...

OpenStudy (anonymous):

left hand side = lny

OpenStudy (amistre64):

\[\frac{f(x)}{(x-a)(x-b)}=\frac{A}{(x-a)}+\frac{B}{(x-b)}\]

OpenStudy (anonymous):

so A/(x+1) +B(x+2)n is the right hand side

OpenStudy (amistre64):

\[\frac{1}{(x+1)(x+2)}=\frac{A}{(x+1)}+\frac{B}{(x+2)}\] \[1=A(x+2)+B(x+1)\] solve A and B for x=-1, and x=-2

OpenStudy (anonymous):

ahhh ok

OpenStudy (amistre64):

1 = A(-1+2) + B(-1+1) ; A = 1 1 = A(-2+2) + B(-2+1) ; B = -1 \[\int\frac{1}{x+1}-\frac{1}{x+2}\]

OpenStudy (anonymous):

ohkay i see, so now

OpenStudy (anonymous):

ln y = ln(x+1) -ln(x+2)

OpenStudy (amistre64):

correct, the base e it all

OpenStudy (anonymous):

oh i forgot to mention y=2 when x=1

OpenStudy (anonymous):

so i find C first dont i?

OpenStudy (anonymous):

ln y = ln(x+1) -ln(x+2) +C

OpenStudy (amistre64):

if you have intial conditions, then sure

OpenStudy (anonymous):

ln2 = ln2-ln3 +C

OpenStudy (anonymous):

C= ln 3

OpenStudy (anonymous):

???

OpenStudy (anonymous):

ln y = ln(x+1) -ln(x+2) + ln3

OpenStudy (anonymous):

y= (x+1) -(x+2) +3

OpenStudy (anonymous):

is that the answer?

OpenStudy (anonymous):

ps i like your picture :)

OpenStudy (amistre64):

\[\Large e^{lna-lnb+c}\] \[\Large e^{ln(a/b)+c}\] \[\Large \frac{a}{b}*c_1\]

OpenStudy (amistre64):

thnx :) \[y = \frac{c(x+1)}{c+2}\] and solve the initial condition

OpenStudy (amistre64):

x+2 under there ... that is

OpenStudy (anonymous):

so y = { (x+1)/(x+2) } multiplied by 3

OpenStudy (amistre64):

that does sound better,

OpenStudy (anonymous):

so \[y= 3(x+1)/x+2\]

OpenStudy (anonymous):

yaaaaayyyyyyyyyyyy :) thank you. i am a fan

OpenStudy (amistre64):

lets step back, and you tell me your initial condition

OpenStudy (anonymous):

i got stuck because of the partial fractions bit. my inital conditions are y=2 when x=1

OpenStudy (anonymous):

where do i put these conditions in?

OpenStudy (amistre64):

\[2 = \frac{c(1+1)}{1+2}\] \[6 = 2c\] yep

OpenStudy (amistre64):

your way was fine :) i was just making sure

OpenStudy (anonymous):

oh so you can also put int the conditions at the end! that's easier i think. because my book is telling me to put in the conditions immediately after integrating. because its a silly book.

OpenStudy (amistre64):

good luck :)

OpenStudy (anonymous):

thank you very much. do u want a medal?

OpenStudy (amistre64):

want? thats not up to me, the thank you is good enough

OpenStudy (anonymous):

thanks very much once again :) you're really helpfull. ill probably be back for more later.

OpenStudy (anonymous):

see you xxxxxxx

OpenStudy (amistre64):

:)